How would you prepare 400 ml of a 0.24 M NaCl solution (MW = 58.44 g/mole) 


1
Expert's answer
2022-06-14T11:35:04-0400

Solution:

Molarity = Moles of solute / Liters of solution

Molarity of NaCl solution = 0.24 M

Liters of NaCl solution = (400 mL) × (1L / 1000 mL) = 0.4 L

Therefore,

Moles of NaCl = Molarity of NaCl solution × Liters of NaCl solution

Moles of NaCl = (0.24 M) × (0.4 L) = 0.096 mol


The molar mass of NaCl is 58.44 g/mol

Therefore,

Mass of NaCl = (0.096 mol NaCl) × (58.44 g NaCl / 1 mol NaCl) = 5.61 g NaCl


Answer: 5.61 g NaCl dissolved in enough water to make 400 mL of solution